A DC voltage generator outputting 10 volts is connected to two capacitors in parallel, who values of 4 and 5 Farads. What is the capacitance? What is the steady-state current?
Answer #1

Page No. Date & CI=4 F C2 =5 F both are in parallel, so Total capacitance (= a+C2 c=4+5 C=9 Farads steady state cursent : - T